OpenStudy (anonymous):

Definite Integrals: This is stumping me! Evaluate the definite integral (to three decimal places): (x-x^2)/(2(cuberoot(x)) dx on [-8,-1]

OpenStudy (anonymous):

Ok, let me re-type this with math type:

OpenStudy (anonymous):

\[\int\limits_{-8}^{-1} (x-x^2)/2\sqrt[3]{x} dx \]

OpenStudy (anonymous):

I just always get nervous when there are factors in the denominator to integrate. I think I should try to bring them up with negative powers, but I'm not entirely sure how to go about that.

OpenStudy (luigi0210):

I think you can separate and integrate separately can't you?

OpenStudy (anonymous):

Integrate separately? Do you mean that I should try to separate the terms? something like this? Int x/2(cuberoot x) - x^2/2(cuberoot x) ? Int (x(2)^-1 (x)^-1/3) - (x^2 (2)^-1 (x)^-1/3) ?

OpenStudy (anonymous):

@TuringTest May I have some help, please? >_< I'm stuck on this one and would love some guidance.

OpenStudy (turingtest):

\[\int\limits_{-8}^{-1} (x-x^2)/2\sqrt[3]{x} dx=\int_{-8}^{-1}\frac12{x-x^2\over x^{1/3}}dx\]can you simplify that?

OpenStudy (anonymous):

\[1/2 \int\limits_{-8}^{-1} x^{2/3} dx - 1/2 \int\limits_{-8}^{-1} x ^{5/3} dx \] like this? o.o

OpenStudy (anonymous):

I pulled the 1/2 out and separates x from x^2 and simplified the exponents, so I have two different terms to integrate. Is this correct?

OpenStudy (turingtest):

you don't have to integrate the two separately, but that is correct

OpenStudy (turingtest):

\[1/2 \int\limits_{-8}^{-1} x^{2/3} dx - 1/2 \int\limits_{-8}^{-1} x ^{5/3} dx=\frac12\int x^{2/3}-x^{5/3}dx\]

OpenStudy (anonymous):

\[(3(8-5x)x^2)/(80\sqrt[3]{x}) \] This is what I ended up getting. I did F(b)-F(a) and got 57.1125 as my answer. It was pretty complicated, but I can type my work if you'd like to look at it.

OpenStudy (turingtest):

it shouldn't be that complicated... I don't know why you reverted back to having the cube root in the denominator

OpenStudy (turingtest):

simply use\[\int x^ndx=\frac{x^{n+1}}{n+1}+C\]

OpenStudy (anonymous):

Hm... let me check. I solved this problem two different ways, and that answer was from my first. I'll try what you just put.

OpenStudy (anonymous):

so 1/2 [ (x^5/3)/5/3 - (x^8/3)/8/3] let me simplify that and then do F(b)-F(a)

OpenStudy (turingtest):

yep :)

OpenStudy (anonymous):

so far I have (-3/10 + 3/16) - (-9.6 + 48)

OpenStudy (anonymous):

I've got -38.513 as my answer. O.o Very off from my first. I think I did this correctly, though. Let me check my work, but it sounds good to me.

OpenStudy (turingtest):

evaluating this integral gives an imaginary answer, so I don't know exactly what they want you to do there's not errors in the problem, right?\[\int_{-8}^{-1}{x-x^2\over 2\sqrt[3]x}dx\]

OpenStudy (anonymous):

No errors that I know of. That is the problem exactly.

OpenStudy (anonymous):

I didn't get an imaginary answer, so regardless, I must have done it incorrectly, huh? As to the imaginary answer, are you sure? I don't think it was a trick question.

OpenStudy (anonymous):

I'm lost now... What should I do, just submit the work I have, try to prove it imaginary, or say that the definite integrable is undeterminable?

OpenStudy (turingtest):

actually wolfram is probably just giving us the imaginary part and we don't need it, let me do it by hand....

OpenStudy (turingtest):

1/2 [ (x^5/3)/5/3 - (x^8/3)/8/3] =\[\frac12[\frac35 x^{5/3}-\frac 38x^{8/3}]|_{-1}^{-8}\]

OpenStudy (anonymous):

Ok thanks, and I'll check over my work. The problem most definitely said DEFINITE integral, but we usually don't work with imaginary numbers in our class, or with calculus in general, right?

OpenStudy (turingtest):

no we shouldn't be dealing with imaginary numbers at this level of calc

OpenStudy (turingtest):

they want the real roots, so in evaluation we want to take \[(-8)^{5/3}\] and the 1/3 in the exponent would cause an imaginary answer if you look at the big picture, but here we only take the real root, so \[(-8)^{5/3}=(-2)^5=-32\]

OpenStudy (anonymous):

Ok I see, so 32 would be acceptable as an answer just dealing with real roots?

OpenStudy (turingtest):

yes, but not as a final answer to the whole integral, I'm just showing you how to evaluate the fractional exponents without getting the imaginary part that wolfram gives

OpenStudy (turingtest):

you still have to plug it all in

OpenStudy (anonymous):

Right, let me try that.

OpenStudy (turingtest):

I'll do the same, let me know what you get

OpenStudy (anonymous):

But for F(b) it would still be the same right? 3(-1)/10 - 3(-1)/16 because it's just one to a power or would the second term be (-1)^8/3 = (-1)^8 = 1 instead? so it is 3(-1)/10 - 3(1)/16 ?

OpenStudy (turingtest):

(-1)^8/3 = 1

OpenStudy (anonymous):

So this is what I got: (-3/10 - 3/16) - (3(-32)/10 - 3(-256)/16) so far

OpenStudy (turingtest):

I think that is right

OpenStudy (turingtest):

wait, (-8)^(8/3) should be positive, since we raise to an even power at the end

OpenStudy (anonymous):

oh so positive 256

OpenStudy (turingtest):

right

OpenStudy (anonymous):

so 57.1125?

OpenStudy (turingtest):

I don't even own a calculator, so let me just write out the simplified form first

OpenStudy (turingtest):

yeah seems right. I always stress about the idea we have made some small arithmetic error somewhere, but I got the same answer, so looks good

OpenStudy (anonymous):

oh ok well I have (-3/10 +3/16) - (-9.6 -48) = -0.4875 + 57.6 = 57.1125 This is how I did it. I use a calculator periodically, kudos to you for not using one! You must be working very hard, so thank you.

OpenStudy (turingtest):

I simplify it and then use google or wolfram, my TI-83 broke :P

OpenStudy (anonymous):

... So 57.113 is what I got the first time I did this. -_- Looks like we came full circle, I guess. XD But I understand the problem much better now, so that's good.

OpenStudy (anonymous):

I'm going to close this question and just look over my work to double-check. Thanks so much for investing so much time into helping me, I really appreciate it! You're getting a medal straightaway!

OpenStudy (turingtest):

Thanks, but my reward is knowing I helped :)